The length of a rectangle is 3 inches more than three times the width. The perimeter is 70 inches. Find the length and width.





Answer by checkley79(3341)

L=3W+3
P=2L+2W
70=2(3W+3)+2W
70=6W+6+2W
70=8W+6
8W=70-6
8W=64
W=64/8
W=8 ANS. FOR THE WIDTH.
L=3*8+3
L=24+3
L=27 ANS. FOR THE LENGTH.
PROOF:
70=2*27+2*8
70=54+16
70=70